OpenStudy (anonymous):

The sum of the height and radius of a cylinder is 6cm. Find the dimensions of the cylineer that will have the largest volume. Find that largest volume (leave answers in terms of pi). The volume of a cylinder is given by V = (pi)(r^2)h.

hartnn (hartnn):

h+r = 6 For largest volume, V' =0 (equating derivative to 0 gives extrema(maxima or minima)) so, what is V' =.... ?

OpenStudy (anonymous):

V' = 2(pi)(r)(h)?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

sorry, for late reply, first use h =6-r and then you get V in terms of 'r' only, only then differentiate.

OpenStudy (stamp):

\[V(r,h)=\pi r^2h\]\[r+h=6\]\[h=6-r\]\[V=\pi r^2(6-r)\] \[V(r)=6\pi r^2-\pi r^3\] \[V'(r)=12\pi r-3\pi r^2\]\[V'(r)=3\pi r(4-r)\]Solve V'(r) = 0 \[0=3\pi r(4-r)\]\[r=0, 4\] We will exclude zero because that's a bogus dimension So r = 4 means h = 2 (recall r + h = 6) (r, h) of (4, 2) gives you your maximum value.

OpenStudy (stamp):

If we had supped r = 6 - h we would have\[V(h)=\pi (6-h)^2h\]\[V(h)=\pi (36-12h+h^2)h\]\[V(h)=36\pi h-12\pi h^2+\pi h^3\]\[V'(h)=36\pi -24\pi h+3\pi h^2\]\[V'(h)=3\pi (12-8h+h^2)\]\[V'(h)=3\pi(h-6)(h-2)\]Solve for V'(h) = 0 h = 6, 2 6 is a bogus dimension because that results in r = 0 so we go with h = 2, and that means h + r = 6 r = 4 Again, (r, h) of (4, 2) gives us our max dimensions.

OpenStudy (stamp):

You see how solving for V(r) instead of V(h) is a cleaner approach